OpenStudy (anonymous):

Solve by factoring or finding square roots (x+5)^2=-36

OpenStudy (anonymous):

NO JUST 36!!!! No NEGATIVE WHATSOEVER!!

OpenStudy (anonymous):

well the way i would do it is like this; (x + 5)^2 = 36 x + 5 = *sqrt(36 x + 5 = 6 x= 6 - 5 x = 1 (x + 5)^2 = 36 (1 + 5)^2 = 36 6^2 = 36 36 = 36 so x = 1

OpenStudy (anonymous):

(x+5)^2=36?

OpenStudy (anonymous):

I think x+5=6 or x+5=-6 x=1 or x=-11

OpenStudy (anonymous):

yeah i agree with what chanh_chung said, it can also be x = -11

OpenStudy (anonymous):

thanks...